### v2

# An inverse problem for semilinear elliptic equations with generalized Kerr-type nonlinearities

Abstract.
We study the inverse problem of reconstructing the shape of unknown inclusions in semilinear elliptic equations with nonanalytic nonlinearities, by extending Ikehata’s enclosure method to accommodate such nonlinear effects. To address the analytical challenges, we construct an approximate solution based on the linearized equation, enabling the enclosure method to operate in this setting. Under suitable structural conditions on the nonlinearity, we establish a shape reconstruction result. The proposed method applies to a broad class of semilinear elliptic equations with non-analytic nonlinearities, including representative examples such as the Kerr-type nonlinearity, which appears in models of nonlinear optics, and the Ginzburg-Landau-type nonlinearity, which models light propagation in nonlinear dissipative media.
Key words and phrases:
enclosure method, semilinear elliptic equation, inverse boundary value problem, Kerr-type nonlinearity, nonlinear optics, Ginzburg-Landau-type nonlinearity, nonlinear dissipative
2020 Mathematics Subject Classification:
35J05, 35J15, 35J61, 35R30
Department of Mathematical Sciences, National Chengchi University, Taipei 116, Taiwan (pzkow@g.nccu.edu.tw). Kow was supported by the National Science and Technology Council of Taiwan, NSTC 112-2115-M-004-004-MY3.
Department of Mathematics, National Cheng Kung University, Tainan 70101, Taiwan (rkuan@ncku.edu.tw). Kuan was supported by the National Science and Technology Council of Taiwan, NSTC 113-2115-M-006-005-MY2.
Acknowledgments
We would like to thank the anonymous readers for their comments and for pointing out several typos in the first draft on arXiv.
1. Introduction
In this paper, we consider the inverse problem of reconstructing an unknown inclusion in a domain from boundary measurements, governed by the following class of semilinear elliptic equations. Let $\Omega\subset\mathbb{R}^{n}$ ($n\geq 2$) be a bounded $C^{1,1}$-domain. We study the problem:
(1.1a) $\Delta u_{f}+q(x,\lvert u_{f}\rvert)u_{f}=0\text{ in $\Omega$},\quad\left.u_{f}\right|_{\partial\Omega}=f.$
[…]
(1.1b) $\lVert q(\cdot,\lvert z_{1}\rvert)z_{1}-q(\cdot,\lvert z_{2}\rvert)z_{2}-q_{0}(\cdot)(z_{1}-z_{2})-q_{1}(\cdot)(\lvert z_{1}\rvert^{\alpha_{1}}z_{1}-\lvert z_{2}\rvert^{\alpha_{1}}z_{2})\rVert_{L^{\infty}(\Omega)}$
$\quad\leq C_{*}\left(\lvert z_{1}\rvert^{\alpha_{2}}+\lvert z_{2}\rvert^{\alpha_{2}}\right)\lvert z_{1}-z_{2}\rvert$
for all $z_{1},z_{2}\in\mathbb{C}$ with $\lvert z_{1}\rvert\leq 1$ and $\lvert z_{2}\rvert\leq 1$. The simplest and typical example of (1.1b) is the well-known Kerr-type nonlinearity, which corresponds to the special case
$q(x,\lvert z\rvert)=q_{0}(x)+q_{1}(x)\lvert z\rvert^{2},$
which arises in models of nonlinear optics, see e.g. [Boy20, MN19] for a physical description. Moreover, this equation arises as the continuum limit of a family of discrete nonlinear Schrödinger equations with long-range lattice interactions, modeling charge transport in biopolymers such as DNA, see e.g. [KLS13]. Using Lemma˜2.2 below, one can readily verify that another example of (1.1b) is the Ginzburg-Landau-type nonlinearity, corresponding to the special case
$q(x,\lvert z\rvert)=q_{0}(x)+q_{1}(x)\lvert z\rvert^{2}+q_{2}(x)\lvert z\rvert^{4},$
which arises in models of light propagation in nonlinear dissipative media, see e.g. [QMM+20]. Furthermore, the nonlinearity considered in [LLST22] is also an instance of the form (1.1b).
Methodology and key difficulties
In this paper, we apply the enclosure method to reconstruct the convex hull of the unknown inclusion. The enclosure method was originally introduced by Ikehata in [Ike99, Ike00], and has since been extended to various types of equations and systems (see, e.g., [BHKS18, BKS15, Ike23, KLS15, NUW11, UW08, WZ13]). The central idea of the method is to substitute a series of special solutions into a designed indicator functional; the resulting values — called the indicator function — are then analyzed. By observing its limiting behavior, one can detect the location of the unknown inclusion.
The method relies on two main components: the special solutions and the indicator functional. In many classical cases, the special solutions are taken to be complex geometrical optics (CGO) solutions. However, in the previous work [Kua24] involving nonlinear equations with power-type nonlinearities, the CGO solutions were replaced by so-called approximate solutions, which were constructed via a Taylor expansion of the associated solution operator. This construction was inspired by the higher-order linearization framework proposed in [LLLS21a, LLLS21b, KU20], and adapted to the setting of the enclosure method.
A main difficulty in our setting is the lack of analyticity (cf. [LLLS21a, LLLS21b, KU20]) of the nonlinear term $z\mapsto q(x,\lvert z\rvert)z$, which renders the Taylor-based construction of approximate solutions in [Kua24] inapplicable. To address this, we introduce a new strategy for constructing approximate solutions suitable for our case. Although this construction differs conceptually from [Kua24], similar types of estimates are still needed to justify the method. These rely on key inequalities developed in [GKM22, KW23], which play a critical role in establishing the required bounds. Further details of this approach are provided in Section˜3.
Mathematical framework and main theorem
To formulate the inverse problem in terms of boundary data, it is essential to ensure that the forward problem is well-posed. In particular, the solution to the boundary value problem must exist uniquely and depend continuously on the boundary data, so that the associated Dirichlet-to-Neumann (DtN) map is well-defined. A detailed discussion of the well-posedness is provided in Section˜2.
We formulate our problem using the following spaces:
$\mathcal{A}^{2}(\Omega):=H^{2}(\Omega)\cap L^{\infty}(\Omega),\quad\mathcal{A}^{\frac{3}{2}}(\partial\Omega):=H^{\frac{3}{2}}(\partial\Omega)\cap L^{\infty}(\partial\Omega)$
each equipped with the norm
$\lVert\cdot\rVert_{\mathcal{A}^{2}(\Omega)}:=\lVert\cdot\rVert_{H^{2}(\Omega)}+\lVert\cdot\rVert_{L^{\infty}(\Omega)},\quad\lVert\cdot\rVert_{\mathcal{A}^{\frac{3}{2}}(\partial\Omega)}:=\lVert\cdot\rVert_{H^{\frac{3}{2}}(\partial\Omega)}+\lVert\cdot\rVert_{L^{\infty}(\partial\Omega)}.$
Both spaces are Banach spaces under these norms. It is well known that $\mathcal{A}^{2}(\Omega)$ and $\mathcal{A}^{\frac{3}{2}}(\partial\Omega)$ are algebras under pointwise multiplication. Specifically, there exists a positive constant $C=C(\Omega)$ such that
(1.2) $\lVert uv\rVert_{\mathcal{A}^{2}(\Omega)}$ $\leq C\lVert u\rVert_{\mathcal{A}^{2}(\Omega)}\lVert v\rVert_{\mathcal{A}^{2}(\Omega)}\quad\text{for all $u,v\in\mathcal{A}^{2}(\Omega)$},$
$\lVert fg\rVert_{\mathcal{A}^{\frac{3}{2}}(\partial\Omega)}$ $\leq C\lVert f\rVert_{\mathcal{A}^{\frac{3}{2}}(\partial\Omega)}\lVert g\rVert_{\mathcal{A}^{\frac{3}{2}}(\partial\Omega)}\quad\text{for all $f,g\in\mathcal{A}^{\frac{3}{2}}(\partial\Omega)$},$
see e.g. [KP88, Kua24] or [Maz11, Section 1.8]. Using Lemma˜2.3 below, one can identify a small parameter
(1.3) $\delta_{0}=\delta_{0}(\Omega,C_{*},\alpha_{1},\alpha_{2},\lVert q_{0}\rVert_{L^{\infty}(\Omega)},\lVert q_{1}\rVert_{L^{\infty}(\Omega)})>0$
such that if
$f\in U_{\delta_{0}}:=\left\{f\in\mathcal{A}^{\frac{3}{2}}(\partial\Omega):\lVert f\rVert_{\mathcal{A}^{\frac{3}{2}}(\partial\Omega)}\leq\delta_{0}\right\},$
then there exists a unique solution $u\in\mathcal{A}^{2}(\Omega)$ of (1.1). Since $\partial\Omega\in C^{1,1}$, using the trace theorem [AH09, Theorem 7.3.11], we can define the Dirichlet-to-Neumann map (DN-map) as
(1.4) $\Lambda_{q}(f):=\partial_{\nu}u_{f}\in H^{1/2}(\partial\Omega)\quad\text{for all $f\in U_{\delta_{0}}$},$
where $\nu$ is the unit outward normal vector on $\partial\Omega$ and $\partial_{\nu}=\nu\cdot\nabla$ is the normal derivative.
Then, with the forward problem properly defined, we move on to the inverse problem studied in this paper: reconstructing an unknown inclusion from boundary measurements in a semilinear setting with non-analytic nonlinearities. We assume that the unknown inclusion is denoted by $D$ satisfying $\overline{D}\subset\Omega$, and represented by a perturbation of the nonlinear coefficient, modeled as
(1.5a) $q_{1}(x)=q_{1,b}(x)+\chi_{D}q_{1,D}(x)\quad\text{for all $x\in\Omega$,}$
where $q_{1,b},q_{1,D}\in L^{\infty}(\Omega)$, $q_{1,b}$ is a known function (representing the “perfect” material), $q_{1,D}$ is unknown (representing a “defect” in the material).
We further assume that the “defect” is detectable in the sense, that is we need a jump condition for the inclusion, that
(1.5b) $\inf_{x\in D}q_{1,D}(x)\geq\mu\quad\text{or}\quad\sup_{x\in D}q_{1,D}(x)\leq-\mu$
for some constant $\mu>0$. The main objective of this paper is to apply a strategy similar to Ikehata’s enclosure method [Ike99, Ike00] to reconstruct the convex hull of the domain $D$ from the boundary measurement $\Lambda_{q}$ defined in (1.4), as stated in the following theorem:
Theorem 1.1 (see Theorem˜3.6 below for a detailed statement).
[…]
It should be noted that our method does not require the exact expression of the nonlinearity $q$. This allows for greater flexibility in the admissible forms of $q$.
Organization
In Section˜2, we establish the well-posedness of the boundary value problem (1.1), followed by the proof of our main result (Theorem˜1.1) in Section˜3.
2. Forward problem
Before we tackle the inverse problem, we first consider the forward problem by establishing the well-posedness of the boundary value problem (1.1). We now introduce the following assumption, see also [GKM22, Assumption 2.1]:
Assumption 2.1 ([KW23, Assumption 1.1]).
There exists $C_{\rm Kerr}>0$, $0\geq q_{0}\in L^{\infty}(\Omega)$ and a $\alpha>0$ such that
$\lVert q(\cdot,\lvert z_{1}\rvert)z_{1}-q(\cdot,\lvert z_{2}\rvert)z_{2}-q_{0}(\cdot)(z_{1}-z_{2})\rVert_{L^{\infty}(\Omega)}\leq C_{\rm Kerr}\left(\lvert z_{1}\rvert^{\alpha}+\lvert z_{2}\rvert^{\alpha}\right)\lvert z_{1}-z_{2}\rvert$
for all $z_{1},z_{2}\in\mathbb{C}$ with $\lvert z_{1}\rvert\leq 1$ and $\lvert z_{2}\rvert\leq 1$.
The following lemma, which demonstrates that ˜2.1 generalizes (1.1b) (with $\alpha=\alpha_{1}$), plays a central role in our proof.
Lemma 2.2 ([GKM22, Lemma A.1]).
The inequality
$\lvert\lvert a\rvert^{\alpha}a-\lvert b\rvert^{\alpha}b\rvert\leq 2(\lvert a\rvert+\lvert b\rvert)^{\alpha}\lvert a-b\rvert$
holds for all $a,b\in\mathbb{C}$ and $\alpha>0$.
Since $0\geq q_{0}\in L^{\infty}(\Omega)$, for each $f\in\mathcal{A}^{\frac{3}{2}}(\partial\Omega)$ one can use Fredholm theory to guarantee that there exists a unique solution $v_{f}\in H^{2}(\Omega)$ to
(2.1) $(\Delta+q_{0})v_{f}=0\text{ in $\Omega$},\quad\left.v_{f}\right|_{\partial\Omega}=f.$
Using [GT01, Theorem 8.30], we know that there exists a unique solution $v_{*}\in H^{1}(\Omega)$ of
(2.2) $\Delta v_{*}=0,\quad\left.v_{*}\right|_{\partial\Omega}=\lVert f\rVert_{L^{\infty}(\partial\Omega)},$
and [GT01, Theorem 8.16] guarantees
(2.3) $\lVert v_{*}\rVert_{L^{\infty}(\Omega)}\leq\lVert f\rVert_{L^{\infty}(\partial\Omega)}.$
The maximum principle for weak solutions [GT01, Theorem 8.19] implies that
(2.4) $v_{*}\geq 0\quad\text{a.e. in $\Omega$.}$
We now see that $v_{*}\pm v_{f}\in H^{1}(\Omega)$ solves
$(\Delta+q_{0})(v_{*}\pm v_{f})=-q_{0}v_{*}\text{ in $\Omega$.}$
Since $(v_{*}\pm v_{f})|_{\partial\Omega}=\lVert f\rVert_{L^{\infty}(\partial\Omega)}\pm f\geq 0$ and $q_{0}\leq 0$ a.e. in $\Omega$, then from (2.4) we see that $q_{0}v_{*}\geq 0$ a.e. in $\Omega$, and again using maximum principle we see that $v_{*}\pm v_{f}\geq 0$ a.e. in $\Omega$, that is,
(2.5) $\lvert v_{f}\rvert\leq v_{*}=\lvert v_{*}\rvert\quad\text{a.e. in $\Omega$.}$
Combining (2.3) and (2.5), we reach
(2.6) $\lVert v_{f}\rVert_{L^{\infty}(\Omega)}\leq\lVert f\rVert_{L^{\infty}(\partial\Omega)}.$
By writing (2.1) as
$\Delta v_{f}=-q_{0}v_{f}\text{ in $\Omega$},\quad\left.v_{f}\right|_{\partial\Omega}=f,$
the trace theorem [AH09, Theorem 7.3.11] (which requires the condition $\partial\Omega\in C^{1,1}$) and the elliptic estimate [GT01, Theorem 8.13] together imply
(2.7) $\lVert v_{f}\rVert_{H^{2}(\Omega)}\leq C\left(\lVert f\rVert_{H^{\frac{3}{2}}(\partial\Omega)}+(1+\lVert q_{0}\rVert_{L^{\infty}(\Omega)})\lVert v_{f}\rVert_{L^{2}(\Omega)}\right).$
Combining (2.7) and (2.6), we conclude
(2.8) $\lVert v_{f}\rVert_{\mathcal{A}^{2}(\Omega)}\leq C(1+\lVert q_{0}\rVert_{L^{\infty}(\Omega)})\lVert f\rVert_{\mathcal{A}^{\frac{3}{2}}(\partial\Omega)}$
for some constant $C=C(\Omega)>1$ which is independent of $q_{0}$.
For each $F\in L^{2}(\Omega)$, again one can use the Lax-Milgram theorem to guarantee that there exists a unique solution $w_{F}\in H_{0}^{1}(\Omega)$ of
(2.9) $(\Delta+q_{0})w_{F}=F\text{ in $\Omega$},\quad\left.w_{F}\right|_{\partial\Omega}=0.$
Since $q_{0}\leq 0$, the Poincaré inequality implies
$\lVert w_{F}\rVert_{H_{0}^{1}(\Omega)}^{2}\leq C\lVert\nabla w_{F}\rVert_{L^{2}(\Omega)}^{2}=C\left(\int_{\Omega}q_{0}\lvert w_{F}\rvert^{2}\,\mathrm{d}x-\int_{\Omega}Fw_{F}\,\mathrm{d}x\right)\leq-C\int_{\Omega}Fw_{F}\,\mathrm{d}x,$
and thus we conclude that
(2.10) $\lVert w_{F}\rVert_{H_{0}^{1}(\Omega)}\leq C\lVert F\rVert_{L^{2}(\Omega)}$
for some constant $C=C(\Omega)>1$ which is independent of $q_{0}$. We now write (2.9) as
$\Delta w_{F}=F-q_{0}w_{F}\text{ in $\Omega$},\quad\left.w_{F}\right|_{\partial\Omega}=0,$
and we use the elliptic estimate [GT01, Theorem 8.13] to see that
$\lVert w_{F}\rVert_{H^{2}(\Omega)}\leq C\left(\lVert F\rVert_{L^{2}(\Omega)}+(1+\lVert q\rVert_{L^{\infty}(\Omega)})\lVert w_{F}\rVert_{L^{2}(\Omega)}\right),$
hence from (2.10) we reach
(2.11) $\lVert w_{F}\rVert_{H^{2}(\Omega)}\leq C(1+\lVert q_{0}\rVert_{L^{\infty}(\Omega)})\lVert F\rVert_{L^{2}(\Omega)}$
for some constant $C=C(\Omega)>1$ which is independent of $q_{0}$. We now further assume that $F\in L^{\infty}(\Omega)$ and let $w_{*}\in H_{0}^{1}(\Omega)$ be the unique solution to
$\Delta w_{*}=\lVert F\rVert_{L^{\infty}(\Omega)}\text{ in $\Omega$},\quad\left.w_{*}\right|_{\partial\Omega}=0.$
By using the maximum principle for weak solutions [GT01, Theorem 8.19] one sees that
(2.12) $w_{*}\geq 0\quad\text{a.e. in $\Omega$.}$
In addition, by using [GT01, Theorem 8.16], there exists a constant $C=C(\Omega)>1$ such that
(2.13) $\lVert w_{*}\rVert_{L^{\infty}(\Omega)}\leq C\lVert F\rVert_{L^{\infty}(\Omega)}.$
We now see that $w_{*}\pm w_{F}\in H_{0}^{1}(\Omega)$ is the unique solution to
$(\Delta+q_{0})(w_{*}\pm w_{F})=\lVert F\rVert_{L^{\infty}(\Omega)}\mp F-q_{0}w_{*}\text{ in $\Omega$},\quad\left.\left(w_{*}\pm w_{F}\right)\right|_{\partial\Omega}=0.$
Since $q_{0}\leq 0$, then from (2.12) we have $\lVert F\rVert_{L^{\infty}(\Omega)}\mp F-q_{0}w_{*}\geq 0$ a.e. in $\Omega$, then by using the maximum principle for weak solutions [GT01, Theorem 8.19] and (2.12) one sees that $w_{*}\pm w_{F}\geq 0$ a.e. in $\Omega$, that is, $\lvert w_{F}\rvert\leq w_{*}=\lvert w_{*}\rvert$ a.e. in $\Omega$, together with (2.13) we now reach
(2.14) $\lVert w_{F}\rVert_{L^{\infty}(\Omega)}\leq C\lVert F\rVert_{L^{\infty}(\Omega)}.$
Now from (2.11) and (2.14) we can define the bounded linear operator
(2.15) $\mathcal{S}:L^{\infty}(\Omega)\rightarrow H_{0}^{1}(\Omega)\cap\mathcal{A}^{2}(\Omega),\quad\mathcal{S}[F]:=w_{F},$
where $w_{F}$ is the unique solution to (2.9) satisfying
(2.16) $\lVert\mathcal{S}[F]\rVert_{\mathcal{A}^{2}(\Omega)}\leq C(1+\lVert q_{0}\rVert_{L^{\infty}(\Omega)})\lVert F\rVert_{L^{\infty}(\Omega)}$
for some constant $C=C(\Omega)>1$ which is independent of $q_{0}$.
We are now ready to prove the well-posedness of (1.1a) under ˜2.1 following the ideas in [KW23, Theorem 2.1], which is based on the contraction mapping theorem.
Lemma 2.3.
Let $\Omega$ be a bounded $C^{1,1}$-domain, let $0\geq q_{0}\in L^{\infty}(\Omega)$ and let $q$ satisfy ˜2.1. There exists a sufficiently small $\delta_{0}=\delta_{0}(\alpha,C_{\rm Kerr},\Omega,\lVert q_{0}\rVert_{L^{\infty}(\Omega)})\in(0,1)$ such that the following statement holds true: if $f\in U_{\delta_{0}}$, then there exists a unique solution $u_{f}\in\mathcal{A}^{2}(\Omega)$ to (1.1a) satisfying
(2.17a) $\lVert u_{f}-v_{f}\rVert_{\mathcal{A}^{2}(\Omega)}\leq C_{\rm Kerr}C^{1+\alpha}(1+\lVert q_{0}\rVert_{L^{\infty}(\Omega)})^{2+\alpha}\lVert f\rVert_{\mathcal{A}^{\frac{3}{2}}(\partial\Omega)}^{1+\alpha}$
and
(2.17b) $\lVert u_{f}\rVert_{\mathcal{A}^{2}(\Omega)}\leq C(1+\lVert q_{0}\rVert_{L^{\infty}(\Omega)})\lVert f\rVert_{\mathcal{A}^{\frac{3}{2}}(\partial\Omega)}$
for some constant $C=C(\Omega)>1$ which is independent of $q_{0},\alpha,C_{\rm Kerr}$, where $v_{f}$ is the unique solution to (2.1).
Proof. Let $f\in\mathcal{A}^{\frac{3}{2}}(\partial\Omega)$ satisfy $\lVert f\rVert_{\mathcal{A}^{\frac{3}{2}}(\partial\Omega)}=\delta\leq\delta_{0}$, where $\delta_{0}>0$ is a small constant to be determined later. From (2.8) we have
(2.18) $\lVert v_{f}\rVert_{\mathcal{A}^{2}(\Omega)}\leq C(1+\lVert q_{0}\rVert_{L^{\infty}(\Omega)})\delta.$
We consider the Banach space
$X_{\delta}:=\left\{w\in H_{0}^{1}(\Omega)\cap\mathcal{A}^{2}(\Omega):\lVert w\rVert_{\mathcal{A}^{2}(\Omega)}\leq\delta\right\},$
and we define
$\mathcal{F}[w]:=-\left(q(\cdot,\lvert w+v_{f}\rvert)-q_{0}\right)(w+v_{f}).$
We first show that
(2.19) $\mathcal{S}[\mathcal{F}[w]]\in X_{\delta}\quad\text{for all $w\in X_{\delta}$}.$
By using (2.16), and consequently using ˜2.1 with the choice $z_{1}=w+v_{f}$ and $z_{2}=0$, we see that
(2.20) $\lVert\mathcal{S}[\mathcal{F}[w]]\rVert_{\mathcal{A}^{2}(\Omega)}\leq C(1+\lVert q_{0}\rVert_{L^{\infty}(\Omega)})\lVert\mathcal{F}[w]\rVert_{L^{\infty}(\Omega)}$
$\quad\leq C_{\rm Kerr}C(1+\lVert q_{0}\rVert_{L^{\infty}(\Omega)})\lVert w+v_{f}\rVert_{L^{\infty}(\Omega)}^{1+\alpha}$
$\quad\leq C_{\rm Kerr}C^{2+\alpha}2^{1+\alpha}(1+\lVert q_{0}\rVert_{L^{\infty}(\Omega)})^{2+\alpha}\delta^{1+\alpha}$
where the last inequality follows from (2.18). If we choose $\delta_{0}>0$ such that
(2.21) $C_{\rm Kerr}C^{2+\alpha}2^{1+\alpha}(1+\lVert q_{0}\rVert_{L^{\infty}(\Omega)})^{2+\alpha}\delta_{0}^{\alpha}\leq 1,$
then we conclude (2.19). Next, we want to show that
(2.22) $\mathcal{S}[\mathcal{F}[\cdot]]$ is a contraction map on $X_{\delta}$.
Let $w_{1},w_{2}\in X_{\delta}$ and compute
$\mathcal{F}[w_{1}]-\mathcal{F}[w_{2}]$ $=q(\cdot,\lvert w_{2}+v_{f}\rvert)(w_{2}+v_{f})-q(\cdot,\lvert w_{1}+v_{f}\rvert)(w_{1}+v_{f})$
$\qquad-q_{0}((w_{2}+v_{f})-(w_{1}+v_{f})).$
Since $\mathcal{S}$ is linear, by using (2.16) and consequently using ˜2.1 with the choice $z_{1}=w_{2}+v_{f}$ and $z_{2}=w_{1}+v_{f}$ we see that
$\lVert\mathcal{S}[\mathcal{F}[w_{1}]]-\mathcal{S}[\mathcal{F}[w_{2}]]\rVert_{\mathcal{A}^{2}(\Omega)}$
$\quad\leq C(1+\lVert q_{0}\rVert_{L^{\infty}(\Omega)})\lVert\mathcal{F}[w_{1}]-\mathcal{F}[w_{2}]\rVert_{L^{\infty}(\Omega)}$
$\quad\leq C_{\rm Kerr}C(1+\lVert q_{0}\rVert_{L^{\infty}(\Omega)})\left(\lVert w_{1}+v_{f}\rVert_{L^{\infty}(\Omega)}^{\alpha}+\lVert w_{2}+v_{f}\rVert_{L^{\infty}(\Omega)}^{\alpha}\right)\lVert w_{1}-w_{2}\rVert_{L^{\infty}(\Omega)}$
$\quad\leq C_{\rm Kerr}C^{1+\alpha}2^{\alpha}(1+\lVert q_{0}\rVert_{L^{\infty}(\Omega)})^{1+\alpha}\delta^{\alpha}\lVert w_{1}-w_{2}\rVert_{L^{\infty}(\Omega)},$
then we conclude (2.22) from (2.21).
From (2.19) and (2.22), the Banach fixed point theorem guarantees that there exists a unique $w_{f}\in X_{\delta}$ such that
(2.23) $w_{f}=\mathcal{S}[\mathcal{F}[w_{f}]].$
We see that $w_{f}$ satisfies
$(\Delta+q_{0})w_{f}=\mathcal{F}[w_{f}]=-\left(q(\cdot,\lvert w_{f}+v_{f}\rvert)-q_{0}\right)(w_{f}+v_{f})\text{ in $\Omega$},\quad w_{f}|_{\partial\Omega}=0.$
We now see that the function $u:=w_{f}+v_{f}\in\mathcal{A}^{2}(\Omega)$ is the unique solution to (1.1a).
On the other hand, from (2.20) and (2.23) we conclude (2.17a). From (2.20) and (2.21) as well as (2.23), we see that
$\lVert w_{f}\rVert_{\mathcal{A}^{2}(\Omega)}\leq\delta=\lVert f\rVert_{\mathcal{A}^{\frac{3}{2}}(\partial\Omega)}.$
Combining the above inequality with (2.8), we conclude (2.17b). ∎
3. Proof of the main result
To clearly present our ideas, we divide the proof of Theorem˜1.1 into several subsections.
3.1. Approximate solutions
In this subsection, we construct approximate solutions suitable for the non-analytic setting. Since the nonlinearity is not analytic, the Taylor-based approach in [Kua24] does not apply. Instead, our construction adopts an iterative perspective: we take the linearized solution $v_{f}$ as the initial approximation and substitute it directly into the nonlinear term, where $v_{f}$ denotes the unique solution of (2.1).
Precisely, letting $\tilde{q}(x,\lvert z\rvert)=q(x,\lvert z\rvert)-q_{0}(x)$ for all $x\in\Omega$ and $z\in\mathbb{C}$, we define $\tilde{u}_{f}$ as the unique solution to the following linear equation:
(3.1) $\Delta\tilde{u}_{f}+q_{0}(x)\tilde{u}_{f}=-\tilde{q}(x,\lvert v_{f}\rvert)v_{f}\text{ in }\Omega,\quad\left.\tilde{u}_{f}\right|_{\partial\Omega}=f.$
We now demonstrate that $\tilde{u}_{f}$ provides a good approximation to $u_{f}$ in the following sense:
Lemma 3.1.
Let $\Omega$ be a bounded $C^{1,1}$-domain, and let $q$ satisfy (1.1b). Then there exists a constant $C=C(\Omega,\lVert q_{0}\rVert_{L^{\infty}(\Omega)},\lVert q_{1}\rVert_{L^{\infty}(\Omega)},\alpha_{1},C_{*})>0$ such that
$\lVert u_{f}-\tilde{u}_{f}\rVert_{\mathcal{A}^{2}(\Omega)}\leq C\lVert f\rVert_{\mathcal{A}^{\frac{3}{2}}(\partial\Omega)}^{1+2\alpha_{1}}\quad\text{for all $f\in U_{\delta_{0}}$,}$
where $\delta_{0}\in(0,1)$ denotes the constant specified in Lemma˜2.3.
Proof. Since $u_{f}-\tilde{u}_{f}=-\mathcal{S}\left[\tilde{q}(\cdot,\lvert u_{f}\rvert)u_{f}-\tilde{q}(\cdot,\lvert v_{f}\rvert)v_{f}\right]$, in other words,
$(\Delta+q_{0})(u_{f}-\tilde{u}_{f})=-\left(\tilde{q}(x,\lvert u_{f}\rvert)u_{f}-\tilde{q}(x,\lvert v_{f}\rvert)v_{f}\right)\text{ in $\Omega$},\quad\left.(u_{f}-\tilde{u}_{f})\right|_{\partial\Omega}=0.$
Using (2.16), we then obtain that
$\lVert u_{f}-\tilde{u}_{f}\rVert_{\mathcal{A}^{2}(\Omega)}\leq C(1+\lVert q_{0}\rVert_{L^{\infty}(\Omega)})\lVert\tilde{q}(\cdot,\lvert u_{f}\rvert)u_{f}-\tilde{q}(\cdot,\lvert v_{f}\rvert)v_{f}\rVert_{L^{\infty}(\Omega)}$
$\quad\leq C(1+\lVert q_{0}\rVert_{L^{\infty}(\Omega)})\left(\begin{aligned} &\lVert\tilde{q}(\cdot,\lvert u_{f}\rvert)u_{f}-\tilde{q}(\cdot,\lvert v_{f}\rvert)v_{f}-q_{1}(\cdot)(\lvert u_{f}\rvert^{\alpha_{1}}u_{f}-\lvert v_{f}\rvert^{\alpha_{1}}v_{f})\rVert\\ &+\lVert q_{1}(\cdot)(\lvert u_{f}\rvert^{\alpha_{1}}u_{f}-\lvert v_{f}\rvert^{\alpha_{1}}v_{f})\rVert_{L^{\infty}(\Omega)}\end{aligned}\right).$
From (2.17b) and (2.18), the smallness of $\delta_{0}\in(0,1)$ ensures that
$\lVert u_{f}\rVert_{L^{\infty}(\Omega)}\leq 1\quad\text{and}\quad\lVert v_{f}\rVert_{L^{\infty}(\Omega)}\leq 1$
which allows us to apply (1.1b). Consequently, we use Lemma˜2.2 and the stability estimates in (2.17) and (2.18), to see that
$\lVert u_{f}-\tilde{u}_{f}\rVert_{\mathcal{A}^{2}(\Omega)}\leq C(1+\lVert q_{0}\rVert_{L^{\infty}(\Omega)})\left(\begin{aligned} &\lVert(\lvert u_{f}\rvert^{\alpha_{2}}+\lvert v_{f}\rvert^{\alpha_{2}})(u_{f}-v_{f})\rVert_{L^{\infty}(\Omega)}\\ &+\lVert q_{1}\rVert_{L^{\infty}(\Omega)}\lVert(\lvert u_{f}\rvert^{\alpha_{1}}+\lvert v_{f}\rvert^{\alpha_{1}})(u_{f}-v_{f})\rVert_{L^{\infty}(\Omega)}\end{aligned}\right)$
$\quad\leq C(1+\lVert q_{0}\rVert_{L^{\infty}(\Omega)})^{3+\alpha}\left(\begin{aligned} &(1+\lVert q_{0}\rVert_{L^{\infty}(\Omega)})^{\alpha_{2}}\lVert f\rVert_{\mathcal{A}^{\frac{3}{2}}(\partial\Omega)}^{\alpha_{2}}\\ &+(1+\lVert q_{0}\rVert_{L^{\infty}(\Omega)})^{\alpha_{1}}\lVert q_{1}\rVert_{L^{\infty}(\Omega)}\lVert f\rVert_{\mathcal{A}^{\frac{3}{2}}(\partial\Omega)}^{\alpha_{1}}\end{aligned}\right)\lVert f\rVert_{\mathcal{A}^{\frac{3}{2}}(\partial\Omega)}^{1+\alpha_{1}}$
$\quad\leq C(\Omega,\lVert q_{0}\rVert_{L^{\infty}(\Omega)},\lVert q_{1}\rVert_{L^{\infty}(\Omega)})\lVert f\rVert_{\mathcal{A}^{\frac{3}{2}}(\partial\Omega)}^{1+2\alpha_{1}},$
which conclude our lemma. ∎
3.2. An indicator functional and its approximation
We define $q_{b}(x,\lvert z\rvert):=q_{0}(x)+q_{1,b}\lvert z\rvert^{\alpha_{1}}$ for all $x\in\Omega$ and $z\in\mathbb{C}$, where $q_{1,b}\in L^{\infty}(\Omega)$ is the function given in (1.5a). By using Lemma˜2.3, there exists a $\delta_{0}\in(0,1)$ such that the following statement holds true: if $f\in U_{\delta_{0}}$, then there exists a unique solution $u_{b,f}\in\mathcal{A}^{2}(\Omega)$ to
$\Delta u_{b,f}+q_{b}(x,\lvert u_{b,f}\rvert)u_{b,f}=0\text{ in $\Omega$},\quad u_{b,f}|_{\partial\Omega}=f.$
Using Lemma˜3.1, we know that there exists a constant $C=C(\Omega,\lVert q_{0}\rVert_{L^{\infty}(\Omega)},\lVert q_{1,b}\rVert_{L^{\infty}(\Omega)},\alpha_{1},C_{*})>0$ such that
$\lVert u_{b,f}-\tilde{u}_{b,f}\rVert_{\mathcal{A}^{2}(\Omega)}\leq C\lVert f\rVert_{\mathcal{A}^{\frac{3}{2}}(\partial\Omega)}^{1+2\alpha_{1}},$
where $\tilde{u}_{b,f}$ is the unique solution to
$\Delta\tilde{u}_{b,f}+q_{0}(x)\tilde{u}_{b,f}=-q_{1,b}(x)\lvert v_{f}\rvert^{\alpha_{1}}v_{f}\text{ in $\Omega$},\quad\left.\tilde{u}_{b,f}\right|_{\partial\Omega}=f,$
where $v_{f}$ is the unique solution to (2.1). Based on the indicator functional in [Kua24] and the structure of our current model, we propose the following variant: Ikehata’s enclosure method [Ike99, Ike00] suggests us to consider the following indicator functional:
$\mathcal{I}(f):=\int_{\partial\Omega}(\partial_{\nu}u_{f}-\partial_{\nu}\tilde{u}_{b,f})\overline{f}\,\mathrm{d}S\quad\text{for all $f\in U_{\delta_{0}}$},$
and the previous paper [Kua24] suggests us to define an auxiliary approximate indicator functional:
$\widetilde{\mathcal{I}}(f):=\int_{\partial\Omega}(\partial_{\nu}\tilde{u}_{f}-\partial_{\nu}\tilde{u}_{b,f})\overline{f}\,\mathrm{d}S\quad\text{for all $f\in U_{\delta_{0}}$.}$
It is important to note that the expression in $\widetilde{\mathcal{I}}$ involves only solutions of linear equations. We now show that $\widetilde{\mathcal{I}}$ is a good approximation of $\mathcal{I}$ in the following sense:
Lemma 3.2.
Let $\Omega$ be a bounded $C^{1,1}$-domain, and let $q$ satisfy (1.1b). Then there exists a constant $C=C(\Omega,\lVert q_{0}\rVert_{L^{\infty}(\Omega)},\lVert q_{1}\rVert_{L^{\infty}(\Omega)},\alpha_{1},C_{*})>0$ such that
$\lvert\mathcal{I}(f)-\widetilde{\mathcal{I}}(f)\rvert\leq C\lVert f\rVert_{\mathcal{A}^{\frac{3}{2}}(\partial\Omega)}^{2+2\alpha_{1}}\quad\text{for all $f\in U_{\delta_{0}}$,}$
where $\delta_{0}\in(0,1)$ is the constant given in Lemma˜2.3.
Proof. Using the trace theorem [AH09, Theorem 7.3.11] (which requires condition $\partial\Omega\in C^{1,1}$), for each $f\in U_{\delta_{0}}$, we compute that
$\lvert\mathcal{I}(f)-\widetilde{\mathcal{I}}(f)\rvert=\left|\int_{\partial\Omega}(\partial_{\nu}u_{f}-\partial_{\nu}\tilde{u}_{f})\overline{f}\,\mathrm{d}S\right|$
$\quad\leq\lVert\partial_{\nu}u_{f}-\partial_{\nu}\tilde{u}_{f}\rVert_{L^{2}(\partial\Omega)}\lVert f\rVert_{L^{2}(\partial\Omega)}$
$\quad\leq C\lVert u_{f}-\tilde{u}_{f}\rVert_{H^{2}(\Omega)}\lVert f\rVert_{L^{2}(\partial\Omega)}.$
Our lemma follows directly from Lemma˜3.1. ∎
We now prove that $\widetilde{\mathcal{I}}$ can be expressed solely in terms of the unique solution $v_{f}$ to (2.1).
Lemma 3.3.
Let $\Omega$ be a bounded $C^{1,1}$-domain, and let $q$ satisfies (1.1b). If we define
$R(x,\lvert z\rvert):=q(x,\lvert z\rvert)-q_{0}(x)-q_{1}(x)\lvert z\rvert^{\alpha_{1}}\quad\text{for all $x\in\Omega$ and for all $z\in\mathbb{C}$,}$
then $\widetilde{\mathcal{I}}$ can be express as
$\widetilde{\mathcal{I}}(f)=-\int_{D}q_{1,D}(x)\lvert v_{f}\rvert^{2+\alpha_{1}}\,\mathrm{d}x-\int_{\Omega}R(x,\lvert v_{f}\rvert)\lvert v_{f}\rvert^{2}\,\mathrm{d}x\quad\text{for all $f\in U_{\delta_{0}}$,}$
where $v_{f}$ is the unique solution to (2.1) and $\delta_{0}\in(0,1)$ denotes the constant specified in Lemma˜2.3.
Remark.
By choosing $z_{1}=v_{f}(x)$ and $z_{2}=0$ in (1.1b), one sees that
(3.2) $\lvert R(x,\lvert v_{f}(x)\rvert)\rvert\leq C_{*}\lvert v_{f}\rvert^{\alpha_{2}}.$
Proof of Lemma˜3.3. We see that the function $z_{f}:=\tilde{u}_{f}-\tilde{u}_{b,f}$ satisfies
$(\Delta+q_{0})z_{f}=-\chi_{D}q_{1,D}(x)\lvert v_{f}\rvert^{\alpha_{1}}v_{f}-R(x,\lvert v_{f}\rvert)v_{f}\text{ in $\Omega$},\quad\left.z_{f}\right|_{\partial\Omega}=0.$
Multiplying the equation above by $\overline{v_{f}}$ and integrating over $\Omega$ yields
$\int_{\Omega}((\Delta+q_{0})z_{f})\overline{v_{f}}\,\mathrm{d}x=-\int_{D}q_{1,D}(x)\lvert v_{f}\rvert^{2+\alpha_{1}}\,\mathrm{d}x-\int_{\Omega}R(x,\lvert v_{f}\rvert)\lvert v_{f}\rvert^{2}\,\mathrm{d}x.$
The lemma then follows by applying integration by parts twice. ∎
3.3. Test data: Complex geometrical optics (CGO) solution
It is important to note that $v_{f}$ is independent of the unknown inclusion $D$, allowing us to select any solution $v_{f}$ to (2.1) with Dirichlet boundary data $f\in U_{\delta_{0}}$. Following Ikehata’s enclosure method [Ike99, Ike00], we choose complex geometrical optics (CGO) solutions, first introduced in [SU87]. In this work, we employ a CGO solution to construct our test data, as described below:
Lemma 3.4.
Let $\Omega$ be a bounded Lipschitz domain and let $\omega,\omega^{\perp}\in\mathbb{S}^{n-1}$ with $\omega\cdot\omega^{\perp}=0$. If $q_{0}\in W^{k,\infty}(\Omega)$ for some $k\in\mathbb{Z}_{\geq 0}$, then for each sufficiently small parameter $h\in(0,1)$ there exists a CGO solution $v$ to $(\Delta+q_{0})v=0$ in $\Omega$, of the form
$v=e^{-\frac{\rho\cdot x}{h}}(1+r_{\rho})$
with $\rho=\omega+\mathbf{i}\omega^{\perp}$ and
(3.3) $\lVert r_{\rho}\rVert_{H^{k}(\Omega)}\leq C_{k}(1+\lVert q_{0}\rVert_{W^{k,\infty}(\Omega)}^{k+1})h$
for some positive constant $C_{k}$ (depends on $k$) which is independent of the parameter $h$.
Remark.
If we choose $k=\lfloor\frac{n}{2}\rfloor+1>\frac{n}{2}$, then the continuous embedding $H^{\lfloor\frac{n}{2}\rfloor+1}(\Omega)\subset\mathcal{A}^{2}(\Omega)$ follows from Sobolev embedding. In this case, from (3.3) we reach
(3.4) $\lVert r_{\rho}\rVert_{\mathcal{A}^{2}(\Omega)}\leq C(1+\lVert q_{0}\rVert_{W^{\lfloor\frac{n}{2}\rfloor+1,\infty}(\Omega)}^{\lfloor\frac{n}{2}\rfloor+2})h$
for some positive constant $C$ which is independent of the parameter $h$.
Proof of Lemma˜3.4. Substituting the ansatz $v=e^{-\frac{\rho\cdot x}{h}}(1+r_{\rho})$ into the equation $(\Delta+q_{0})v=0$ in $\Omega$, we obtain
$0$ $=h^{2}e^{\frac{\rho\cdot x}{h}}(\Delta+q_{0})\left(e^{-\frac{\rho\cdot x}{h}}(1+r_{\rho})\right)$
$=h^{2}e^{\frac{\rho\cdot x}{h}}(\Delta+q_{0})(e^{-\frac{\rho\cdot x}{h}}r_{\rho})+h^{2}q_{0}.$
In terms of the differential operator
$\mathcal{P}_{h}:=h^{2}e^{\frac{\rho\cdot x}{h}}(\Delta+q_{0})(e^{-\frac{\rho\cdot x}{h}}\cdot)=h^{2}\left(\Delta-\frac{2}{h}\rho\cdot\nabla+q_{0}\right),$
we see that
(3.5) $\mathcal{P}_{h}r_{\rho}=h^{2}e^{\frac{\rho\cdot x}{h}}(\Delta+q_{0})(e^{-\frac{\rho\cdot x}{h}}r_{\rho})=-h^{2}q_{0}.$
Step 1: Existence of $r_{\rho}$. We first prove that there exists a solution $r_{\rho}\in H^{1}(\Omega)$ to (3.5). Since $q_{0}\in W^{0,\infty}(\Omega)=L^{\infty}(\Omega)$ and $\Omega$ is bounded Lipschitz, by using [DSFKSU07, Proposition 2.4] (which also can be found in [Sal07, Proposition 5.2]), there exists $\tilde{r}\in H^{1}(\Omega)$ such that
$e^{\frac{\omega\cdot x}{h}}h^{2}(\Delta+q_{0})(e^{-\frac{\omega\cdot x}{h}}\tilde{r})=-h^{2}e^{-\frac{i\omega^{\perp}\cdot x}{h}}q_{0}\quad\text{in $\Omega$}$
with
$h\lVert\tilde{r}\rVert_{L^{2}(\Omega)}+h^{2}\lVert\nabla\tilde{r}\rVert_{L^{2}(\Omega)}\leq Ch^{2}\lVert q_{0}\rVert_{L^{2}(\Omega)}$
for some positive constant $C$ which is independent of $h$ and $\tilde{r}$. We see that $r_{\rho}:=e^{\mathbf{i}\frac{\omega^{\perp}\cdot x}{h}}\tilde{r}\in L^{2}(\Omega)$ satisfies (3.5) and
$\lVert r_{\rho}\rVert_{L^{2}(\Omega)}\leq Ch\lVert q_{0}\rVert_{L^{2}(\Omega)}$
which proves (3.3) with $k=0$.
Step 2: Regularity of $r_{\rho}$. Having already established (3.3) for $k=0$, it remains to prove it for $k=m+1$, assuming the inductive hypothesis that (3.3) holds for $k\leq m$ and $r_{\rho}\in H^{m+1}(\Omega)$. Let $\boldsymbol{\beta}$ be a multiindex with $\lvert\boldsymbol{\beta}\rvert=m+1$, we obtain formally
[…]
$\quad=h^{2}\left(-\frac{2}{h}\rho\cdot\nabla(\partial^{\boldsymbol{\beta}}r_{\rho})+\Delta(\partial^{\boldsymbol{\beta}}r_{\rho})+\sum_{\boldsymbol{\gamma}\leq\boldsymbol{\beta}}(\partial^{\boldsymbol{\gamma}}q_{0})(\partial^{\boldsymbol{\beta}-\boldsymbol{\gamma}}r_{\rho})\right)$
$\quad=\mathcal{P}_{h}(\partial^{\boldsymbol{\beta}}r_{\rho})+h^{2}\sum_{\boldsymbol{\gamma}\leq\boldsymbol{\beta},\boldsymbol{\gamma}\neq 0}(\partial^{\boldsymbol{\gamma}}q_{0})(\partial^{\boldsymbol{\beta}-\boldsymbol{\gamma}}r_{\rho}).$
The induction hypothesis asserts that
$\mathcal{P}_{h}(\partial^{\boldsymbol{\beta}}r_{\rho})=-h^{2}\left(\partial^{\boldsymbol{\beta}}q_{0}+\sum_{\boldsymbol{\gamma}\leq\boldsymbol{\beta},\boldsymbol{\gamma}\neq 0}(\partial^{\boldsymbol{\gamma}}q_{0})(\partial^{\boldsymbol{\beta}-\boldsymbol{\gamma}}r_{\rho})\right)\in L^{2}(\Omega).$
If we write $\tilde{r}_{\boldsymbol{\beta}}=e^{-\mathbf{i}\frac{\omega^{\perp}\cdot x}{h}}\partial^{\boldsymbol{\beta}}r_{\rho}$, then we reach
$h^{2}e^{\frac{\omega\cdot x}{h}}(\Delta+q_{0})(e^{-\frac{\omega\cdot x}{h}}\tilde{r}_{\boldsymbol{\beta}})=-h^{2}e^{-\mathbf{i}\frac{\omega^{\perp}\cdot x}{h}}\left(\partial^{\boldsymbol{\beta}}q_{0}+\sum_{\boldsymbol{\gamma}\leq\boldsymbol{\beta},\boldsymbol{\gamma}\neq 0}(\partial^{\boldsymbol{\gamma}}q_{0})(\partial^{\boldsymbol{\beta}-\boldsymbol{\gamma}}r_{\rho})\right).$
We apply the Carleman estimate in [DSFKSU07, Proposition 2.4] to observe that
$h\lVert\tilde{r}_{\boldsymbol{\beta}}\rVert_{L^{2}(\Omega)}+h^{2}\lVert\nabla\tilde{r}_{\boldsymbol{\beta}}\rVert_{L^{2}(\Omega)}$
$\quad\leq Ch^{2}\left(\lVert q_{0}\rVert_{W^{m+1,\infty}(\Omega)}+\sum_{j=1}^{m+1}\lVert q_{0}\rVert_{W^{j,\infty}(\Omega)}\lVert r_{\rho}\rVert_{H^{m+1-j}}\right)$
$\quad\leq Ch^{2}(1+\lVert q_{0}\rVert_{W^{m+1,\infty}}^{m+2}),$
which conclude (3.3) for $k=m+1$. ∎
We slightly modify the CGO solution in order to construct test data whose Dirichlet trace belongs to class $U_{\delta_{0}}$.
Lemma 3.5.
Let $\Omega$ be a bounded $C^{1,1}$-domain, let $\omega,\omega^{\perp}\in\mathbb{S}^{n-1}$ with $\omega\cdot\omega^{\perp}=0$ and let $q_{0}\in W^{\lfloor\frac{n}{2}\rfloor+1,\infty}(\Omega)$. Given a parameter $t\in\mathbb{R}$ and a large parameter $J>0$, we consider the CGO solution
$v_{h}=e^{-\frac{J-t}{h}}e^{-\frac{\rho\cdot x}{h}}(1+r_{\rho})=e^{-\frac{J}{h}}e^{-\frac{1}{h}(\omega\cdot x-t+\mathbf{i}\omega^{\perp}\cdot x)}(1+r_{\rho})$
and we denote its Dirichlet trace $f_{h}:=v_{h}|_{\partial\Omega}$. If we write
$b(\omega):=\inf_{x\in\Omega}\omega\cdot x,$
then
$\lVert f_{h}\rVert_{\mathcal{A}^{\frac{3}{2}}(\partial\Omega)}\leq Ch^{-2}e^{-\frac{J}{h}}e^{\frac{t-b(\omega)}{h}}\quad\text{for all sufficient small $h\in(0,1)$}$
for some positive constant $C$ which is independent of the parameter $h$.
Proof. For each $j,k=1,\cdots,n$, one computes
$\partial_{j}\partial_{k}v_{h}=e^{-\frac{J}{h}}e^{-\frac{1}{h}(\rho\cdot x-t)}\left(\left(\frac{\rho_{j}}{h}\right)\left(\frac{\rho_{k}}{h}\right)(1+r_{\rho})-\left(\frac{\rho_{k}}{h}\right)\partial_{j}r_{\rho}-\left(\frac{\rho_{j}}{h}\right)\partial_{k}r_{\rho}+\partial_{j}\partial_{k}r_{\rho}\right),$
and then using (3.4) we see that
$\lVert\partial_{j}\partial_{k}v_{h}\rVert_{L^{2}(\Omega)}\leq Ce^{-\frac{J}{h}}e^{\frac{t-b(\omega)}{h}}h^{-2}$
for some positive constant $C$ which is independent of $h$. Since
$\lVert v_{h}\rVert_{L^{2}(\Omega)}\leq Ce^{-\frac{J}{h}}e^{\frac{t-b(\omega)}{h}},$
then the trace theorem [AH09, Theorem 7.3.11] (which requires the condition $\partial\Omega\in C^{1,1}$) implies
$\lVert f_{h}\rVert_{\mathcal{A}^{\frac{3}{2}}(\partial\Omega)}\leq C\lVert v_{h}\rVert_{H^{2}(\Omega)}\leq Ce^{-\frac{J}{h}}e^{\frac{t-b(\omega)}{h}}h^{-2}$
for some positive constant $C$ which is independent of $h$. On the other hand, one sees that
$\lVert f_{h}\rVert_{L^{\infty}(\partial\Omega)}\leq 2\sup_{x\in\partial\Omega}\lvert e^{-\frac{J}{h}}e^{-\frac{1}{h}(\omega\cdot x-t+\mathbf{i}\omega^{\perp}\cdot x)}\rvert\leq 2e^{-\frac{J}{h}}e^{\frac{t-b(\omega)}{h}}$
thus proving the lemma. ∎
As an immediate consequence, if $J>B(\omega)-b(\omega)$ and $t<B(\omega)$, where
$B(\omega):=\sup_{x\in\Omega}\omega\cdot x,$
then for all sufficiently small $h\in(0,1)$, we have $f_{h}\in U_{\delta_{0}}$. This ensures the CGO solution Lemma˜3.5 can be substituted into the indicator functionals $\mathcal{I}$ and $\widetilde{\mathcal{I}}$.
3.4. Conclusion
We can now finally prove our main theorem as follows:
Theorem 3.6 (see also Theorem˜1.1 above).
Let $\Omega\subset\mathbb{R}^{n}$ ($n\geq 2$) be a bounded $C^{1,1}$-domain. Assume there exist $0\geq q_{0}\in W^{\lfloor\frac{n}{2}\rfloor+1,\infty}(\Omega)$, $q_{1}\in L^{\infty}(\Omega)$ with (1.5), a constant $C_{*}>0$, an integer $\alpha_{1}\in\mathbb{Z}_{\geq 2}$, a real number $\alpha_{2}>\alpha_{1}$, such that the nonlinear term $q$ satisfies (1.1b). Let $t_{*}:=\inf_{x\in D}\omega\cdot x$ and let $f_{h}$ be the function given in Lemma˜3.5. For each $\omega\in\mathbb{S}^{n-1}$, we choose
(3.6) $J>\max\left\{\frac{(B(\omega)-b(\omega))(2+2\alpha_{1})}{\alpha_{1}},\frac{(B(\omega)-b(\omega))(2+\alpha_{2})}{\alpha_{2}-\alpha_{1}}\right\},\quad b(\omega)<t<B(\omega).$
Then the following statements hold:
  1. (i)
    If $\{\omega\cdot x\leq t\}\cap\overline{D}=\emptyset$, that is, $t<t_{*}$, then
    $\lim_{h\searrow 0}e^{\frac{(\alpha_{1}+2)J}{h}}\lvert\mathcal{I}(f_{h})\rvert=0.$
  2. (ii)
    If $\{\omega\cdot x\leq t\}\cap D\neq\emptyset$, that is, $t>t_{*}$, then
    $\lim_{h\searrow 0}e^{\frac{(\alpha_{1}+2)J}{h}}\lvert\mathcal{I}(f_{h})\rvert$$\infty$.
Remark.
Theorem˜3.6 implies that for any fixed direction $\omega\in\mathbb{S}^{n-1}$, one can determine whether the hyperplane $\omega\cdot x=t$} intersects the inclusion $D$ by testing the limiting behavior of $e^{\frac{(\alpha_{1}+2)J}{h}}\lvert\mathcal{I}(f_{h})\rvert$ as $h\to 0$. By varying $t$, we can identify the critical value $t_{*}$ such that the hyperplane $\omega\cdot x=t_{*}$} just touches $D$, although the limiting behavior at $t=t_{*}$ itself is not known. By applying this procedure in every direction $\omega\in\mathbb{S}^{n-1}$, we reconstruct the convex hull of $D$ and thereby conclude Theorem˜1.1.
Let us first make some observations before proving Theorem˜3.6. Using Lemma˜3.3, we insert the CGO solution $v_{h}$ given in Lemma˜3.5 into $\widetilde{\mathcal{I}}$ to see that
$\widetilde{\mathcal{I}}(f_{h})=-\int_{D}q_{1,D}(x)\lvert v_{h}\rvert^{2+\alpha_{1}}\,\mathrm{d}x-\int_{\Omega}R(x,\lvert v_{h}\rvert)\lvert v_{h}\rvert^{2}\,\mathrm{d}x.$
It follows that
$e^{\frac{(\alpha_{1}+2)J}{h}}\mathcal{I}(f_{h})$ $=-e^{\frac{(\alpha_{1}+2)J}{h}}\int_{D}q_{1,D}(x)\lvert v_{h}\rvert^{2+\alpha_{1}}\,\mathrm{d}x-e^{\frac{(\alpha_{1}+2)J}{h}}\int_{\Omega}R(x,\lvert v_{h}\rvert)\lvert v_{h}\rvert^{2}\,\mathrm{d}x$
$\quad+e^{\frac{(\alpha_{1}+2)J}{h}}(\mathcal{I}(f_{h})-\widetilde{\mathcal{I}}(f_{h})).$
By using Lemma˜3.2, Lemma˜3.5 and (3.6), we see that
$e^{\frac{(\alpha_{1}+2)J}{h}}\lvert\mathcal{I}(f_{h})-\widetilde{\mathcal{I}}(f_{h})\rvert\leq Ce^{\frac{(\alpha_{1}+2)J}{h}}\lVert f_{h}\rVert_{\mathcal{A}^{\frac{3}{2}}(\partial\Omega)}^{2+2\alpha_{1}}$
$\quad\leq Ch^{-4-4\alpha_{1}}e^{\frac{(\alpha_{1}+2)J}{h}}e^{-\frac{J(2+2\alpha_{1})}{h}}e^{\frac{(t-b(\omega))(2+2\alpha_{1})}{h}}$
$\quad=Ch^{-4-4\alpha_{1}}\exp\left(-\frac{1}{h}(\alpha_{1}J-(t-b(\omega))(2+2\alpha_{1}))\right)$
$\quad\leq Ch^{-4-4\alpha_{1}}\exp\left(-\frac{1}{h}(\alpha_{1}J-(B(\omega)-b(\omega))(2+2\alpha_{1}))\right)$
$\quad\rightarrow 0\quad\text{as $h\searrow 0$.}$
By using (3.2), (3.4) and (3.6), we see that
$e^{\frac{(\alpha_{1}+2)J}{h}}\left|\int_{\Omega}R(x,\lvert v_{h}\rvert)\lvert v_{h}\rvert^{2}\,\mathrm{d}x\right|\leq C_{*}e^{\frac{(\alpha_{1}+2)J}{h}}\int_{\Omega}\lvert v_{h}\rvert^{2+\alpha_{2}}\,\mathrm{d}x$
$\quad\leq C_{*}e^{\frac{(\alpha_{1}+2)J}{h}}(1+\lVert r_{\rho}\rVert_{L^{\infty}(\Omega)})^{2+\alpha_{2}}\int_{\Omega}e^{-\frac{J}{h}(2+\alpha_{2})}e^{-\frac{2+\alpha_{2}}{h}(\omega\cdot x-t)}\,\mathrm{d}x$
$\quad\leq C\left(1+h+\lVert q_{0}\rVert_{W^{\lfloor\frac{n}{2}\rfloor+1,\infty}(\Omega)}^{\lfloor\frac{n}{2}\rfloor+2}h\right)^{2+\alpha_{2}}e^{-\frac{J}{h}(\alpha_{2}-\alpha_{1})}e^{-\frac{2+\alpha_{2}}{h}(b(\omega)-B(\omega))}$
$\quad\rightarrow 0\quad\text{as $h\searrow 0$.}$
Therefore we see that (provided that the limit exist)
(3.7) $\lim_{h\searrow 0}e^{\frac{(\alpha_{1}+2)J}{h}}\lvert\mathcal{I}(f_{h})\rvert=\lim_{h\searrow 0}e^{\frac{(\alpha_{1}+2)J}{h}}\left|\int_{D}q_{1,D}(x)\lvert v_{h}\rvert^{2+\alpha_{1}}\,\mathrm{d}x\right|.$
We now ready to prove our main result Theorem˜3.6.
Proof of Theorem˜3.6(i). Using Lemma˜3.5 and (3.4) we estimate
$e^{\frac{(\alpha_{1}+2)J}{h}}\left|\int_{D}q_{1,D}(x)\lvert v_{h}\rvert^{2+\alpha_{1}}\,\mathrm{d}x\right|$
$\quad\leq e^{\frac{(\alpha_{1}+2)J}{h}}\lVert q_{1,D}\rVert_{L^{\infty}(D)}(1+\lVert r_{\rho}\rVert_{L^{\infty}(\Omega)})^{2+\alpha_{1}}\left|\int_{D}e^{-\frac{J}{h}(2+\alpha_{1})}e^{-\frac{2+\alpha_{1}}{h}(\omega\cdot x-t)}\,\mathrm{d}x\right|$
$\quad\leq C\lVert q_{1,D}\rVert_{L^{\infty}(D)}\left(1+h+\lVert q_{0}\rVert_{W^{\lfloor\frac{n}{2}\rfloor+1,\infty}(\Omega)}^{\lfloor\frac{n}{2}\rfloor+2}h\right)^{2+\alpha_{2}}e^{-\frac{2+\alpha_{1}}{h}(t_{*}-t)}$
$\quad\rightarrow 0\quad\text{as $h\searrow 0$,}$
since $t<t_{*}$, hence, in conjunction with (3.7), proving the result. ∎
Proof of Theorem˜3.6(ii). Using (1.5b), we obtain
$e^{\frac{(\alpha_{1}+2)J}{h}}\left|\int_{D}q_{1,D}(x)\lvert v_{h}\rvert^{2+\alpha_{1}}\,\mathrm{d}x\right|\geq e^{\frac{(\alpha_{1}+2)J}{h}}\mu\int_{D}\lvert v_{h}\rvert^{2+\alpha_{1}}\,\mathrm{d}x$
$\quad\geq\mu\int_{D}e^{-\frac{2+\alpha_{1}}{h}(\omega\cdot x-t)}\lvert 1+r_{\rho}\rvert^{2+\alpha_{1}}\,\mathrm{d}x.$
Note that (3.4) ensures that
$\lVert r_{\rho}\rVert_{L^{\infty}(\Omega)}\leq\left(1+\lVert q_{0}\rVert_{W^{\lfloor\frac{n}{2}\rfloor+1,\infty}(\Omega)}^{\lfloor\frac{n}{2}\rfloor+2}\right)h\leq\frac{1}{2}\quad\text{for all sufficiently small $h\in(0,1)$,}$
therefore we reach
$e^{\frac{(\alpha_{1}+2)J}{h}}\left|\int_{D}q_{1,D}(x)\lvert v_{h}\rvert^{2+\alpha_{1}}\,\mathrm{d}x\right|\geq 2^{-2-\alpha_{1}}\mu\int_{D}e^{-\frac{2+\alpha_{1}}{h}(\omega\cdot x-t)}\,\mathrm{d}x.$
Since $D$ is open, then there exists $\epsilon>0$ such that $t_{*}<t-\epsilon$ and the set $\{\omega\cdot x\leq t-\epsilon\}\cap D$ has positive measure. We now have
$e^{\frac{(\alpha_{1}+2)J}{h}}\left|\int_{D}q_{1,D}(x)\lvert v_{h}\rvert^{2+\alpha_{1}}\,\mathrm{d}x\right|\geq 2^{-2-\alpha_{1}}\mu\int_{\{\omega\cdot x\leq t-\epsilon\}\cap D}e^{-\frac{2+\alpha_{1}}{h}(\omega\cdot x-t)}\,\mathrm{d}x$
$\quad\geq 2^{-2-\alpha_{1}}\mu\int_{\{\omega\cdot x\leq t-\epsilon\}\cap D}e^{\frac{2+\alpha_{1}}{h}\epsilon}\,\mathrm{d}x=2^{-2-\alpha_{1}}\mu\lvert\{\omega\cdot x\leq t-\epsilon\}\cap D\rvert e^{\frac{2+\alpha_{1}}{h}\epsilon}$
$\quad\rightarrow+\infty\quad\text{as $h\searrow 0$,}$
hence, in conjunction with (3.7), proving the result. ∎